OpenStudy (anonymous):

3+10+17+24+...+7003+7010, How do I find the "n' for the formula Sn = n/2 (a1+an)?

OpenStudy (campbell_st):

well I'd use the equation for a term in an arithmetic sequence rather than the sum equation. you know the nth term is 7010 and the 1st term is 3 as well as the common difference is 7 so substitute them into \[a_{n}= a_{1} + (n - 1) \times d\] which becomes \[7010 = 3 + (n - 1) \times 7\] now solve for n

OpenStudy (anonymous):

Thanks!